OpenStudy (anonymous):

What is the exponent of e in the 4th term of the binomial expansion? (e^2 - 4f)^5 a)6 b)2 c)4 d)3

OpenStudy (mimi_x3):

Well, you can use this to find the 4th term in the expansion or the pascal traingle I suppose.. \[\sum_{r=0}^{5}*\left(\begin{matrix}5 \\ r\end{matrix}\right)*(e^2)^{5-r}*(-4f)^{r}\] \[\sum_{r=0}^{5}*\left(\begin{matrix}5 \\ r\end{matrix}\right)*(e^2)^{5-r}*(-4)^{r}*(f)^{r}\]

OpenStudy (mimi_x3):

\[\sum_{r=0}^{5}*\left(\begin{matrix}5 \\ r\end{matrix}\right)*e^{10-2r}*(-4)^{r}*(f)^{r}\]

OpenStudy (mimi_x3):

If I remember correctly; the 4th term is when r=3

OpenStudy (anonymous):

so whats the exponent on the e?

OpenStudy (shubhamsrg):

4th term can either be for r=3 or r=2..

OpenStudy (mimi_x3):

woops, on the exponent of \(e\) then it's still \(r=3\) right?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

anybody understand? sorry im so lost

OpenStudy (mimi_x3):

maybe @ParthKohli: have an easier method; this is the only method which I remember

Parth (parthkohli):

Yeah, I do have one

OpenStudy (mimi_x3):

Well, my method is easy; than the pascals triangle..

Parth (parthkohli):

If you take the expansion of \(( a + b)^2 \) for example — you see something. \(\Rightarrow a^2 + 2ab + b^2\) In the first term, we have 2 as the exponent of a and 0 as the exponent of b. In the second term, we have 1 - 1. In the third term, we have 0 - 2. The exponents always adds up to n.

Parth (parthkohli):

So you just start with: 5 - 0 <--- first term 4 - 1<---- second term 3 - 2<---- third term 2 - 3<----fourth term

Parth (parthkohli):

Now \(e\) would look something like \((e^2)^2 = e^4\) right?

Parth (parthkohli):

@Mimi_x3 What was I called for? lol. To explain this method or not? :P

OpenStudy (anonymous):

yep thanks!!!

Parth (parthkohli):

You're welcome.

OpenStudy (mimi_x3):

To conclude the method that I used: \[\binom{5}{3} *(e)^{10-2(3)} *(-4)^{3} *f^{3} => -640e^{4}*f^{3}\] which is more convenient if you want to find the whole term. (: